OpenStudy (anonymous):

What is the ordered pair for the equation y = x + 3 if x = 1?

OpenStudy (diyadiya):

y=x+3 x=1 y=1+3 y= cn you do?

OpenStudy (diyadiya):

and ordered pair will be like this (x,y) =(1,-- )

OpenStudy (anonymous):

1,3?

OpenStudy (diyadiya):

No y=1+3 what is 1+3 ?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

so 1,4?

OpenStudy (diyadiya):

Yeah (1,4) is the ordered pair